A 25 ft ladder is placed against a vertical wall. The foot of the ladder is 7 feet from the base of the wall. If the foot of the ladder slips 8 feet , then the top of the ladder will slide
A 4 ft b 6 ft c 8ft d 20 ft e none of these

Answer by richwmiller(17219)

25^2=7^2+x^2
625=49+x^2
x=24
625=15^2+y^2
y=20
24-20=4
the top of the ladder slips 4 ft